OpenStudy (anonymous):

Look at this function: x y −12 −8 −8 −6 −4 −4 0 −2 Mina said that the function is linear. Jenna said that the function is nonlinear. Which of the following explains who is correct? Mina, because for every 2-unit increase in x, there is an increase in y by 4 Mina, because for every 4-unit increase in x, there is an increase in y by 2 Jenna, because the point (0, −2) does not lie on the straight line that contains the other points Jenna, because the point (−12, −8) does not lie on the straight line that contains the other points

OpenStudy (anonymous):

sorry its all smooshed together

OpenStudy (freckles):

well let's look at the reasons... what number is x increasing by each time? what number is y increasing by each time?

OpenStudy (anonymous):

x=4

OpenStudy (anonymous):

y=2

OpenStudy (freckles):

right because -12+4 gives us the next number -8 for x and yes -8+2 gives us the next number -6 for y so we definitely know it isn't option A because that explanation is kinda backwards

OpenStudy (freckles):

so B looks like a good candidate we can also look at Jenna's reasonings

OpenStudy (anonymous):

yeah ok cross out A

OpenStudy (freckles):

linear functions can be written as f(x)=mx+b

OpenStudy (freckles):

anything that can be written as f(x)+mx+b is a liner function m and b can both be zero

OpenStudy (anonymous):

hmm so C

OpenStudy (freckles):

oops f(x)=mx+b * do you know how to attempt to find this line given two points say (-4,-4) and (0,-2)

OpenStudy (anonymous):

or is that crossed out

OpenStudy (anonymous):

i son't sorry but i can try figureing it out:()

OpenStudy (freckles):

the slope represents m \[m=\frac{y_1-y_2}{x_1-x_2} =\frac{-2-(-4)}{0-(-4)}\]

OpenStudy (freckles):

what is -2+4 and 0+4?

OpenStudy (anonymous):

2 and 4

OpenStudy (freckles):

right! \[f(x)=\frac{2}{4}x+b \\ \text{ we can reduce the slope number there } \\ f(x)=\frac{1}{2}x+b \\ \text{ now using either point we can come up with b } \\ f(-4)=\frac{1}{2}(-4)+b=-4 \\ \\ \text{ so we need \to solve } \\ \frac{1}{2}(-4)+b=-4 \\ -2+b=4\] do you know what b is?

OpenStudy (anonymous):

some glitching is happening sorry

OpenStudy (anonymous):

didnt mean to say that... i get it

OpenStudy (anonymous):

B= 6?

OpenStudy (freckles):

well I left the - off on my 4 in the next line the last should have been -2+b=-4

OpenStudy (freckles):

so b=-2

OpenStudy (anonymous):

oh ok thanks

OpenStudy (freckles):

\[f(x)=\frac{1}{2}x-2 \\ \] let's look at your points and see this linear equation works for all of them: \[f(-12)=\frac{1}{2}(-12)-2=-6-2=-8 \text{ \check mark for the first point } \\ f(-8)=\frac{1}{2}(-8)-2=-4-2=-6 \text{ \check for mark the second point } \\ f(-4)=\frac{1}{2}(-4)-2=-2-2=-4 \text{ \check mark for the third point }\] and you check the last one if you want to anyways what this says is jenna is totally wrong since we were able to describe the points in the chart with a linear function

OpenStudy (freckles):

so if you can spot a similar pattern like mina did you can determine it is linear from that

OpenStudy (anonymous):

so neither is mina and not A either so its B?

OpenStudy (freckles):

yep it is B it is the correct answer with a correct reasoning

OpenStudy (anonymous):

thx now i know what i am doing you have helped a lot thank you so much:)

OpenStudy (freckles):

np